The average number of dining chairs assembled per five-working day week was 450 in an assembly process of a furniture plant. 6 workers were working in this process throughout the measurement interval. What is the labor productivity in this process?

  • a.15 chairs/worker/day

  • b.20 chairs/worker/day

  • c.90 chairs/worker/day

  • d.60 chairs/worker/day

  • f.75 chairs/worker/day
Answer #1

The average number of chairs assembled per 5 working day week by 6 workers = 450 chairs

The number of chairs assembled in 1 working day by 6 workers = 450/5 = 90 chairs

as 90 chairs are assembled by 6 workers in 1 working day, this cannot be the answer.

6 workers were working in the process during the measurement interval which is 1 week (5 days)

6 workers finished 90 chairs in a day.

1 worker would finish 90 /6 = 15 chairs a day

Hence the labor productivity = 15 chairs/worker/day

Another way :

The average number of chairs assembled per 5 working day week by 6 workers = 450 chairs

On an average 1 worker assembles 450 / 6 = 75 chairs in 5 working day week.

As 75 chairs are assembled by 1 worker in 5 days, this is not correct answer

1 worker assembles an average of 75 / 5 = 15 chairs in 1 day

Hence the labor productivity = 15 chairs/worker/day

Options b and d are also not correct as these are unrelated numbers.

The correct answer is a. 15 chairs / worker / day.